2 blocks each of lenght x, width x+3, and hight is x, are cut of a block of length 3x+4, width x+3, hieght x. all dimensions are given in centimeter. Find the volume of resulting solid?
Answer by nerdybill(7384)

The volume of the resulting solid (in terms of 'x') is:
"vol of resulting solid?" = "vol of original block" - "vol of the 2 cut blocks"
"vol of resulting solid?" = x(x+3)(3x+4) - 2(x)(x)(x+3)
"vol of resulting solid?" = (x^2+3x)(3x+4) - (2x^2)(x+3)
"vol of resulting solid?" = (3x^3+4x^2+9x^2+12x) - (2x^3+6x^2)
"vol of resulting solid?" = (3x^3+13x^2+12x) - (2x^3+6x^2)
"vol of resulting solid?" = (3x^3+13x^2+12x) - 2x^3-6x^2
"vol of resulting solid?" = (x^3+13x^2+12x) -6x^2
"vol of resulting solid?" = x^3+7x^2+12x